Please solve. Thanks!
1.)y=(x+3)^2
2.)y=(x-4)^2

Answer by checkley77(12844)

y=(x+3)^2
y=x^2+6x+9
------------------
y=(x-4)^2
y=x^2-8x+16 setting these 2 equation equal we get:
-----------------
x^2+6x+9=x^2-8x+16
x^2-x^2+6x+8x=16-9
14x=7
x=7/14
x=1/2 answer.
y=(1/2+3)^2
y=3.5^2
y=12.25 answer.
y=(1/2-4)^2
y=-3.5^2
y=12.25 answer.